OpenStudy (anonymous):

best answer gets medal and a fan The table shows the results of drawing colored marbles from a bag. Based on theoretical probability, which is the best prediction for the number of times a blue marble would be drawn in 1000 trials? A. 150 B. 200 C. 250 D. 300

OpenStudy (anonymous):

here is the chart

OpenStudy (anonymous):

can u help mee

OpenStudy (anonymous):

Mathmale can u figre it out plz

OpenStudy (anonymous):

can anyone figure it out

OpenStudy (anonymous):

dude math male can u help me

OpenStudy (mathmale):

First, I'm assuming that you must have some experience with observed probability, either in class or in online materials. Is that correct? Supposing that you had a sack of 20 marbles, some blue, some yellow, some red, and through actual counting find that the number of blue marbles is 3. What is the probability that if you start with this sack of 20 marbles, that you'll draw a blue one?

OpenStudy (mathmale):

Another question for you: Please look at the table and determine how many marbles, total, are in the bag described.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 40 marbels actully

OpenStudy (mathmale):

and how many of them are blue?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

@mathmale r u still gonna help or no

OpenStudy (anonymous):

i thout u left

OpenStudy (mathmale):

Good. Therefore, the probability that you'll draw a blue marble is what? Write the appropriate fraction and then reduce it.

OpenStudy (anonymous):

10/40=1/4

OpenStudy (anonymous):

umm i dont know

OpenStudy (anonymous):

whats the formula

OpenStudy (anonymous):

hey@mathmale thx big time for helping me

OpenStudy (mathmale):

So, the theoretical prob. that you'll draw a blue marble from the sack of 40 is 1/4. Now, let's apply that. supposing you keep drawing marbles from the bag and always replacing whichever marble you drew. You do this 1000 times. (Sounds like fun!) How many blue marbles would you be likely to draw Iin 1000 trials?

OpenStudy (mathmale):

I changed my last post because I'd made a mistake there. The correct version is above.

OpenStudy (anonymous):

so is the formula 10/1000=10%

OpenStudy (mathmale):

If the P is 1/4 and the number of trials is 1000, how many blue marb. are you likely to draw, total, always replacing the drawn marbles?

OpenStudy (mathmale):

Would you mind explaining how you got that? What's your reasoning?

OpenStudy (anonymous):

i got 1/4 by doing 10/40=1/4

OpenStudy (mathmale):

that's fine: that's the P of drawing a blue marble each time you draw one from the bag of 40 marbles. But the question is different now. suppose you have nothing else to do and so you sit on your front port and draw marbles 1000 times. how many are likely to be blue?

OpenStudy (anonymous):

hey i need help

OpenStudy (mathmale):

I was waiting for you to try the math. If the theoretical prob. of drawing a blue marb. is 1/4, and you draw and repl. a marb. 1000 times, then the expected number of blue marbs. that you'd draw would be (1/4) (1000) = ?

OpenStudy (anonymous):

i really just need answer plz but if u cant then teach me plz

OpenStudy (anonymous):

250 @mathmale

OpenStudy (anonymous):

@Darth_Beast dont just ask for the answer try and figure out what mathmale is saying